OpenStudy (anonymous):

Write an equivalent exponential or logarithmic equation. e2 = 6x

hartnn (hartnn):

take log on both sides, then you can use the log property for left side \(\Large \log a^b = b \log a\)

OpenStudy (anonymous):

so itll be loge^2=6 logx ???